please help me solve this. its linear systems by substitution
3x+2y=5
5x-9y=-4

Answer by mananth(16949)

3 x + 2 y = 5 .............1

5 x -9 y = -4 .............2
Eliminate y
multiply (1)by 9
Multiply (2) by 2
27 x 18 y = 45
10 x -18 y = -8
Add the two equations
37 x = 37
/ 37
x = 1
plug value of x in (1)
3 x + 2 y = 5
3 + 2 y = 5
2 y = 5 -3
2 y = 2
y = 1
x=1, y=1